# The impact of medium-width bands on the selection, and subsequent luminosity function measurements, of high-z galaxies

### v1

@@ 4.2 Results when repeating the experiment on a simulated catalogue
Our selection criteria is designed to identify sources at , however there are a large number of sources around this boundary which scatter above and below this selection limit as a consequence of the use of photo-z’s. To simplify discussion of completeness and contamination, we consider sources with a true redshift of in the following subsections. We also break up this discussion into two redshift bins; and . In the lower redshift bin, the presence of the Balmer break and rest-optical emission lines may result in different behaviours when applying medium-width bands when compared to higher redshifts where the available UV emission lines are typically weaker and photometry generally probes the UV slope. We present the results of this process in Figure 7 and discuss them in more detail in the following subsections.
@@ 4.2.1 High-z Sample Completeness
Beginning with the redshift bin of , we find that completeness (Number of sources at selected divided by the total number of sources) is generally high when considering sources at . Without any medium-band information, this completeness is over 80% at the mark, while including even a single medium band can substantially raise this to 90%. As would likely be expected, increasing the number of medium-width bands leads to an increase in completeness, though the largest benefit comes from the inclusion of a single medium-width band and subsequent bands are less impactful. This improvement with a single medium-width band is noted in the previous subsection, where the biggest jump in the JOF sample overlap was when only F410M was added to the catalogues. From a sample completeness perspective, these results thus support the general ‘deep-field’ design of including a single medium-width band and F410M presents one of the better balances of sensitivity and longest wavelength (to study e.g. emissions lines to highest redshifts or the Balmer break up to ) as has been employed by many early survey designs as a balance of cost verses effectiveness.
Examining higher redshifts at , completeness is very high () for sources detected to , however the number of these sources present in the simulation a these luminosities is very low. As observed for the sources, completeness turns over and falls off below . The use of 0, 1, 2, or even 7 medium-width bands does not provide a significant difference in the completeness curve, due to the lack of any strong, defining features present in the rest frame UV. It is only in our hypothetical scenarios that completeness is improved, with either the inclusion of the reddest medium bands or greater depths in the blue broad bands leading to notable improvements in the 5-8 regime.
@@ 4.2.2 High-z Sample Contamination
Contamination is when members of an unwanted sub-population enter a sample selection. In our case, contaminant galaxies are those which are of significantly lower redshift () than the target sample of . The contaminating fraction is defined as the number of these low-z sources divided by the total number of sources (contaminant plus true) selected. In all of our simulated runs, we find that contamination is generally low for the JADES Origins Field. Even without the use of medium-width bands, contamination is zero for sources detected to and begins to rise at increasing rates as the level is approached. The primary reason is with such high significance in the detections, the measured depth of the Lyman break discounts the potential Balmer break alternative in the bluer bands.
@@
Examining contamination in the two primary redshift bins we explore, contamination rates are much higher for the highest redshift galaxies at compared to . These contamination rates at very high redshifts may be an overestimation, since recent JWST findings have identified greater numbers of galaxies than first expected. The UV luminosity density measured from recent studies of the UV LF (e.g. Adams et al., 2024; Donnan et al., 2024) is around a factor of 10 higher than JAGUAR estimations at (Williams et al., 2018) and this discrepancy increases to higher redshift. As a result, the relative numbers of contaminant galaxies and real galaxies may differ compared to reality. The emphasis here is that contamination becomes a concerning risk at the regime and that cuts could produce more conservative/robust samples (Supporting the methods of e.g. McLeod et al., 2024).
@@ 5.2 Changes in Photometric Redshifts & Selections for High-z Sources
We begin our investigation by assessing which objects meet our full selection criteria after all photometric bands are degraded by the same level. The initial selection criteria finds 50 sources at , this decreases to 46, 21, 12 when the images are degraded by the three steps described above. The number of consistent objects between the samples generated from degraded imaging was 34, 17, 9, meaning that with each degradation step, sources were both lost and gained from the sample selection. For the remainder of this section we explore what these sources are and why their classification as robust high-z galaxies (or not) changed.
@@
The vast majority of sources which are selected as high-z sources in shallow catalogues but found to be confidently low-z in deeper catalogues are within the 5-8.5 regime in the respective shallow catalogue. Using the numbers of contaminants found, under the assumption that the photo-z using the full depth imaging is correct, the contamination rate in the 5-8.5 regime ranges from 20-38% between catalogues. There is a single instance of a strongly detected source switching from low-z in deep catalogues to high-z in shallower catalogues. This object is the spectroscopically confirmed source (Carniani et al., 2024). It is found that the increasing background results in this source’s successful selection as the faint extended emission from its neighbour is lost to noise, creating the expected non-detections from a Lyman dropout.
@@ 6 Measuring the faint-end of the UV Luminosity Function with the JOF
The results of this process are presented in Figure 10f. The deep imaging provided by the JOF field enables galaxies to be identified down to at and at . Compared to Adams et al. (2024), which compiled 6 JWST surveys together over 180 square arcminutes of area, the JOF field expands the faint end coverage by 0.5mags. We conduct our parametric fits to the UV LF using these new JOF results, the Adams et al. (2024) results and, where possible, constraints from the ground-based UltraVISTA survey from Donnan et al. (2023). Our observations are consistent with the high number density of galaxies found by other JWST surveys that have pushed the depth frontier at , including the NGDEEP (Leung et al., 2023) and MIDIS surveys (Pérez-González et al., 2023).
-20.25 | 7.03 | 6.52 |
-19.25 | 14.24 | 10.44 |
-18.25 | 85.30 | 40.00 |
-18.25 | 50.60 | 24.27 |
-18.75 | 19.07 | 12.20 |
@@
9 | Sch/LM | – | |||
9 | Sch/MCMC | – | |||
9 | DPL/LM | ||||
9 | DPL/MCMC | ||||
10.5 | Sch/LM | – | |||
10.5 | Sch/MCMC | – | |||
10.5 | DPL/LM | ||||
10.5 | DPL/MCMC | ||||
10.5 | Sch/LM | – | |||
10.5 | Sch/MCMC | – | |||
10.5 | DPL/LM | ||||
10.5 | DPL/MCMC | ||||
12.5 | Sch/LM | – | |||
12.5 | Sch/MCMC | – | |||
12.5 | DPL/LM | ||||
12.5 | DPL/MCMC |
@@ 6.1 Fitting the UV LF
To parameterise the UV LF and describe any potential evolution, it has been common process to fit the UV LF of galaxies with either a Schechter (SCH: power law with an exponential decline Schechter, 1976) or Double Power Law (DPL) parameterisation. These both have a normalisation parameter (), a characteristic luminosity/magnitude where the shape changes () and a faint-end slope (). The DPL functional form has an additional bright-end slope () in replacement to the exponential decay in the Schechter form.
@@
Relative to results from Adams et al. (2024), uncertainties in the faint-end slopes (when left free) have decreased by around 25-30% thanks to the increased depth provided by the JOF field allowing us to identify intrinsically fainter galaxies (mags).
@@ 6.2 A consistent faint-end of the UV LF between photometric catalogues
We find that for the and UV LF, the majority of galaxies used are consistent between the C-ALL, C-0, C-1 and C-2 catalogues, though up to 20% of the galaxies occupying the faintest bin can change. These galaxies fall into two categories; 1) they either have a photo-z close to the redshift bin boundary which changes slightly between catalogues (e.g. in the bin) scattering them in/out or 2) they are among the faintest in the sample and are no longer selected due to changes in confidence relative to low-z solutions in different catalogues. For the redshift bin, the C-ALL catalogue employs the use of 3 galaxies, while C-2 and C-1 use 1 galaxy and C-0 uses no galaxies. The scatter in redshift is larger in this regime, with two galaxies in C-ALL now or in C-0. Additionally, completeness and contamination are also sensitive. For C-ALL, the average completeness is 83% for the galaxies used while it drops to 40% for C-2 and C-1, with this factor of two in completeness making up some of the difference in measured number density obtained from using 1 or 3 galaxies.
@@
When using a different aperture size, we also observe minor changes to the measured UV LF that match expectations. The deeper coverage provided by the use of smaller apertures enables intrinsically fainter galaxies (0.25-0.5 mags) to be identified while the brighter galaxies found in the wider aperture catalogue have the majority recovered with similar redshifts and properties. We provide the final UV LF fits to the 0.2as version of the UV LF on account of its constraining power on the faint-end slope in Table 4 and provide the points in Table 3.
@@ 6.3 The UV Luminosity Density
When fitting a trivial log-linear evolution (e.g. Donnan et al., 2023; McLeod et al., 2024; Donnan et al., 2024) to our updated UV luminosity density measurements (including the result from Adams et al. (2024), we find a slope of and normalisation of . The steepness of this evolution is shallower than initially measured in Adams et al. (2024), closer to the value measured by McLeod et al. (2024) but steeper than found in Donnan et al. (2024). When compared to theoretical predictions, there is a general bimodality in the star formation rate density expected, with a selection of simulations and analytic prescriptions predicting low star formation rates (SCSAM Yung et al. (2024), Thesan Kannan et al. (2022), UniverseMachine Behroozi et al. (2019) in addition to Harikane et al. (2022)) and a selection predicting higher (FLARES Wilkins et al. (2023), Delphi Dayal et al. (2022), Ferrara et al. (2023) in addition to Behroozi & Silk (2015)). At , our observations more closely match to those predicting lower SFRD, while at the opposite is the case, with observations more closely matching predictions of high SFRD. Further increased volume and sightlines probing fainter luminosities (e.g. the complete JADES survey and lensing clusters from e.g. PEARLS, CANUCS and GLIMPSE) will ultimately be required in order fully constrain the shape of the UV LF at and .
@@ 7 Conclusions
Our fiducial measurement of the faint end of the UV LF at redshifts 8.5 to 13.5 agree with previous estimates from early JWST data, evidencing high numbers of moderately faint galaxies () at this early time. The estimated star formation rate density of the Universe at is higher than previously measured in Adams et al. (2024) and closer to other early JWST-based measurements.
@@
It is also worth stressing that development continues to improve high-z selections. This work displays the expected reliability and issues facing currently employed systems. New templates of galaxy SEDs are continuously being tested (e.g. Larson et al., 2023; Clausen et al., 2024), understanding of IGM attenuation is improving (e.g. Asada et al., 2024) and the application of informed priors can be used to more conservatively examine fainter sources (e.g. Donnan et al., 2024). Increasing spectroscopic samples of populations which can masquerade as high-z galaxies (be it galactic brown dwarfs, dusty line emitters or Balmer break galaxies) or objects occupying unexplored parameter spaces (e.g. RUBIES de Graaff et al., 2024) will be pivotal to refining high redshift selections and SED template sets. With large spectroscopic programmes like JADES, and now with CAPERS underway, a more statistical analysis of SED modelling comparing to real spectra will soon be enabled.

### v2

@@ 4.2 Results when repeating the experiment on a simulated catalogue
Our selection criteria is designed to identify sources at , however there are a large number of sources around this boundary which scatter above and below this selection limit as a consequence of the use of photo-z’s. To simplify discussion of completeness and contamination, we consider sources with a true redshift of in the following subsections. We also break up this discussion into two redshift bins; and . In the lower redshift bin, the presence of the Balmer break and rest-optical emission lines in the reddest bands may result in different behaviours when adding medium-width bands when compared to higher redshifts, where the available UV emission lines are typically weaker and photometry generally probes the UV slope. We present the results of this process in Figure 8 and discuss them in more detail in the following subsections.
@@ 4.2.1 High-z Sample Completeness
Beginning with the redshift bin of , we find that completeness (Number of sources at selected divided by the total number of sources) is generally high when considering sources at . Without any medium-band information, this completeness is over 80% at the mark, while including even a single medium band can substantially raise this to 90%. As would likely be expected, increasing the number of medium-width bands leads to an increase in completeness, though the largest benefit comes from the inclusion of a single medium-width band and subsequent bands are less impactful. This improvement with a single medium-width band is noted in the previous subsection, where the biggest jump in the JOF sample overlap was when only F410M was added to the catalogues.
To better understand the benefits that the single F410M medium band provides, we explore the reasons why completeness increases by in the bin when it is included. Examining the full sample of 215 of these sources, we find that all have high-z primary solutions, but the dominant population are those with true redshifts of but have a photo-z of , resulting in their non-selection with our criteria. These sources have strong O[III]/H emission boosting F444W fluxes. Without medium band information, the strong emission lines are sometimes placed at the bluest end of the F444W filter providing a photo-z of . The addition of F410M reveals an absence of emission lines at the bluest side of F444W and the F300M to F335M colour also provides a more precise Balmer break position. Such sources that have systematically underestimated photo-z’s account for of the sources contained within this completeness increase. For the remaining sources they fail our constraints when no medium-width bands are available. From a sample completeness perspective, these results thus support the general ‘deep-field’ design of including a single medium-width band and F410M presents one of the better balances of sensitivity and longest wavelength (to study e.g. rest-optical emissions lines or the Balmer break up to ) as has been employed by many early survey designs as a balance of cost verses effectiveness.
Examining higher redshifts at , completeness is very high () for sources detected to , however the number of these sources present in the simulation a these luminosities is very low. As observed for the sources, completeness turns over and falls off below . The use of 0, 1, 2, or even 7 medium-width bands does not provide a significant difference in the completeness curve, due to the lack of any strong, defining features present in the rest frame UV. Medium bands primarily aid in increasing or decreasing the constraint relative to the next-best low-z solution. It is only in our hypothetical scenarios that completeness is improved, with either the inclusion of the reddest medium bands (providing some constraints on the Balmer break at ) or greater depths in the blue broad bands leading to notable improvements in the 5-8 regime.
@@ 4.2.2 High-z Sample Contamination
Contamination is when members of an unwanted sub-population enter a sample selection. In our case, contaminant galaxies are those which are of significantly lower redshift () than the target sample of . The contaminating fraction is defined as the number of these low-z sources divided by the total number of sources (contaminant plus true) selected. It should be noted that JAGUAR was produced in the pre-JWST era. Comparing its intrinsic UV LF to observations across cosmic history (e.g. Bouwens et al., 2021), JAGUAR number densities typically are slightly lower but within a factor of . However, at very high redshifts beyond previous HST limits (), JAGUAR number densities are around factor of 20 below new observations (e.g. Adams et al., 2024; Donnan et al., 2024; Whitler et al., 2025). Consequently, the ratio of potential low-z contaminants and high-z galaxies differs from reality and a JAGUAR-derived contamination rate would be too high. To compensate, correctly identified galaxies at are up-weighted by a factor of 10 in our contamination calculations.
In all of our simulated runs, we find that contamination is generally low for the JADES Origins Field. Even without the use of medium-width bands, contamination is zero for sources detected to and begins to rise at increasing rates as the level is approached. The primary reason is with such high significance in the detections, the measured depth of the Lyman break discounts the potential Balmer break alternative in the bluer bands.
@@ 5.2 Changes in Photometric Redshifts & Selections for High-z Sources
We begin our investigation by assessing which objects meet our full selection criteria after all photometric bands are degraded by the same level. We identify 46, 21, 12 sources when the images are degraded by the three steps described above. The number of consistent objects between the samples generated from degraded imaging was 34, 17, 9, meaning that with each degradation step, sources were both lost and gained from the sample selection. For the remainder of this section we explore what these sources are and why their classification as robust high-z galaxies (or not) changed.
@@
The vast majority of sources which are selected as high-z sources in shallow catalogues but found to be confidently low-z in deeper catalogues are within the 5-8.5 regime in the respective shallow catalogue. Using the numbers of contaminants found, under the assumption that the photo-z using the full depth imaging is correct, the contamination rate in the 5-8.5 regime ranges from 20-38% between catalogues. There is a single instance of a strongly detected source switching from low-z in deep catalogues to high-z in shallower catalogues. This object is JADES-GS-z14-0 (Carniani et al., 2024; Schouws et al., 2025; Carniani et al., 2025). It is found that the increasing background results in this source’s successful selection as the faint extended emission from its neighbour is lost to noise, creating the expected non-detections from a Lyman dropout.
@@ 6 Measuring the faint-end of the UV Luminosity Function with the JOF
The results of this process are presented in Figure 11f. The deep imaging provided by the JOF field enables galaxies to be identified down to at and at . Compared to Adams et al. (2024), which compiled 6 JWST surveys together over 180 square arcminutes of area, the JOF field expands the faint end coverage by 1mag. We conduct our parametric fits to the UV LF using these new JOF results, the Adams et al. (2024) results and, where possible, constraints from the ground-based UltraVISTA survey from Donnan et al. (2023). Our observations are consistent with the high number density of galaxies found by other JWST surveys that have pushed the depth frontier at , including the NGDEEP (Leung et al., 2023) and MIDIS surveys (Pérez-González et al., 2023) and the JADES teams own UV LF measurements (Robertson et al., 2024; Whitler et al., 2025).
| -20.00 | 14.28 | 10.85 |
| -19.00 | 58.40 | 28.52 |
| -18.00 | 101.2 | 47.9 |
| -19.25 | 45.33 | 39.87 |
| -18.25 | 50.60 | 29.38 |
| -19.00 | 14.30 | 10.50 |
@@
| 9 | Sch/LM | – | |||
| 9 | Sch/MCMC | – | |||
| 9 | DPL/LM | ||||
| 9 | DPL/MCMC | ||||
| 10.5 | Sch/LM | – | |||
| 10.5 | Sch/MCMC | – | |||
| 10.5 | DPL/LM | ||||
| 10.5 | DPL/MCMC | ||||
| 10.5 | Sch/LM | – | |||
| 10.5 | Sch/MCMC | – | |||
| 10.5 | DPL/LM | ||||
| 10.5 | DPL/MCMC | ||||
| 12.5 | Sch/LM | – | |||
| 12.5 | Sch/MCMC | – | |||
| 12.5 | DPL/LM | ||||
| 12.5 | DPL/MCMC |
@@ 6.1 Fitting the UV LF
To parameterise the UV LF and describe any potential evolution, it has been common process to fit the UV LF of galaxies with either a Schechter (SCH: power law with an exponential decline Schechter, 1976) or Double Power Law (DPL) parameterisation. These both have a normalisation parameter (), a characteristic luminosity/magnitude where the shape changes () and a faint-end slope (). The DPL functional form has an additional bright-end slope () in replacement to the exponential decay in the Schechter form, adding more flexibility when describing brighter galaxies.
@@
Relative to results from Adams et al. (2024), uncertainties in the faint-end slopes (when left free) have decreased by around % thanks to the increased depth provided by the JOF field allowing us to identify intrinsically fainter galaxies (mag).
@@ 6.2 A consistent faint-end of the UV LF between photometric catalogues
We find that for the and UV LF, the majority of galaxies used are consistent between the C-ALL, C-0, C-1 and C-2 catalogues, though up to 20% of the galaxies occupying the faintest bin can change. These galaxies fall into two categories; 1) they either have a photo-z close to the redshift bin boundary which changes slightly between catalogues (e.g. in the bin) scattering them in/out or 2) they are among the faintest in the sample and are no longer selected due to changes in confidence relative to low-z solutions in different catalogues. For the redshift bin, the C-ALL catalogue employs the use of 2 galaxies, while C-2 and C-1 use 1 galaxy and C-0 uses no galaxies. The scatter in redshift is larger in this regime, with two galaxies in C-ALL now or in C-0. Additionally, completeness and contamination are also sensitive. For C-ALL, the average completeness is 83% for the galaxies used while it drops to 40% for C-2 and C-1, with this factor of two in completeness making up some of the difference in measured number density obtained from using 1 or 2 galaxies.
@@
When using a different aperture size, we also observe minor changes to the measured UV LF that match expectations. The deeper coverage provided by the use of smaller apertures enables intrinsically fainter galaxies ( mags) to be identified while the brighter galaxies found in the wider aperture catalogue have the majority recovered with similar redshifts and properties. We provide the final UV LF fits to the 0.2as version of the UV LF on account of its constraining power on the faint-end slope in Table 4 and provide the points in Table 3.
6.2.1 Testing the impact of Damped Lyman Alpha systems
Most photo-z studies in the early Universe make use of the Lyman break formed by the attenuation of UV flux from the neutral IGM (Madau, 1995; Inoue et al., 2014). At very high redshifts, the Lyman break is essentially a step function. However, NIRSpec observations have shown a prevalence of damped Lyman alpha (DLA) systems in the early Universe. DLA’s arise from neutral gas in close proximity to a galaxy and can soften the shape of the Lyman break. This can result in systematic offsets in photometric redshifts as reported in Asada et al. (2024). We experiment with the implementation of the DLA attenuation described in Asada et al. (2024) in our photo-z process. Since the JAGUAR SAM doesn’t include such attenuation, we focus on the application to the real JOF catalogues.
We find that including DLA features in our template fitting results in systematic shifts towards lower-z for our sample. The magnitude of these shifts is an average of 0.08 in redshift between and 0.18 between (). Using our fiducial catalogue, we find that the UV LF does not change significantly when using these slightly modified redshifts due ot the low numbers and broad bins. These redshift biases of up to 0.18 at will likely be more impactful when working with larger, more statistical samples of galaxies from e.g. multi-survey studies.
6.3 Comparisons to Other Observational Studies
In this subsection we examine our binned number densities and best fitting Schechter/DPL parameters with those of other JWST-based observations undertaken in the past couple of years. At , our new number densities at are in close agreement to observations from most other observational studies (Finkelstein et al., 2024; Leung et al., 2023; Pérez-González et al., 2023; Donnan et al., 2023, 2024; Adams et al., 2024; Whitler et al., 2025), showing a general consensus across multiple different survey fields. The only exception is the GLASS field which may be slightly overdense at the bright end (Castellano et al., 2022).
At , the data points from Adams et al. (2024) are on the lower-end of the number densities measured from multiple JWST-based studies though are generally consistent within with the exception of the point at which is over below the works of Donnan et al. (2024), Leung et al. (2023) and Pérez-González et al. (2023). On the other hand, the fainter data points added by the inclusion of the JOF at and are in very close agreement to the observations from those same studies. This could indicate a potential overestimation of sample completeness at the faint-end of Adams et al. (2024), though it agrees with pre-JWST observations from Oesch et al. (2018), or it could be driven by the impacts of cosmic variance. Most studies reaching this regime primarily make use of ultra-deep blank fields (NGDEEP, MIDIS, JOF etc.) which are all located the wider GOODS-S region. The inclusion of more ultra-deep or gravitationally lensed fields will be required to fully assess the impact of cosmic variance at the very faint end.
At , our observations again agree with the other surveys reaching , now including Robertson et al. (2024), with the same caveat that these are all located within the wider GOODS-S field. Previous fits within Adams et al. (2024) resulted in a relatively large uncertainty in the normalisation which has now been further constrained to the upper 50%, showing that the UV LF evolves relatively slowly from to . However, our fits provide galaxy number densities that are around (or dex) higher than the multi-survey study of Donnan et al. (2024), further raising the question of potential cosmic variance induced systematics which will require more independent deep field programmes to solve.
@@ 6.4 The UV Luminosity Density
When fitting a trivial log-linear evolution (e.g. Donnan et al., 2023; McLeod et al., 2024; Donnan et al., 2024) to our updated UV luminosity density measurements (including the result from Adams et al. (2024), we find a slope of and normalisation of . The steepness of this evolution is shallower than initially measured in Adams et al. (2024), almost identical to the value measured by McLeod et al. (2024) but steeper than found in Donnan et al. (2024). When compared to theoretical predictions, there is a general bimodality in the star formation rate density expected, with a selection of simulations and analytic prescriptions predicting low star formation rates (SCSAM Yung et al. (2024), Thesan Kannan et al. (2022), UniverseMachine Behroozi et al. (2019) in addition to Harikane et al. (2022)) and a selection predicting higher (FLARES Wilkins et al. (2023), Delphi Dayal et al. (2022), Ferrara et al. (2023) in addition to Behroozi & Silk (2015)). At , our observations more closely match to those predicting lower SFRD, while at the opposite is the case, with observations more closely matching predictions of high SFRD. Further increased volume and sightlines probing fainter luminosities (e.g. the complete JADES survey and lensing clusters from e.g. PEARLS, CANUCS and GLIMPSE) will ultimately be required in order fully constrain the shape of the UV LF at and .
@@ 7 Conclusions
Our fiducial measurement of the faint end of the UV LF at redshifts 8.5 to 13.5 agree with previous estimates from early JWST data, evidencing high numbers of moderately faint galaxies () at this early time. The estimated star formation rate density of the Universe at is higher than previously measured in Adams et al. (2024) and closer to other early JWST-based measurements. However, there is the caveat of cosmic variance, as all other surveys reaching these depths have been concentrated within the wider GOODS-South region (e.g. NGDEEP and MIDIS Leung et al., 2023; Pérez-González et al., 2023).
@@
It is also worth stressing that development continues to improve high-z selections. This work displays the expected reliability and issues facing currently employed systems. New templates of galaxy SEDs are continuously being tested (e.g. Larson et al., 2023; Clausen et al., 2024), understanding of IGM attenuation is improving (e.g. Asada et al., 2024) and the application of informed priors can be used to more conservatively examine fainter sources (e.g. Donnan et al., 2024). Increasing spectroscopic samples of populations which can masquerade as high-z galaxies (be it galactic brown dwarfs, dusty line emitters or Balmer break galaxies) or objects occupying unexplored parameter spaces (e.g. RUBIES de Graaff et al., 2024) will be pivotal to refining high redshift selections and SED template sets. With large spectroscopic programmes like JADES, and now with CAPERS underway, a more statistical analysis of SED modelling comparing to real spectra will soon be enabled. CAPERS in particular will explore high-z sources from multiple, independently produced input catalogues which will be key to revealing why different groups select different sources and reveal more insights into the properties of contaminant sources (see e.g. Castellano et al., 2025).